OpenStudy (anonymous):

The equation y = 18x shows that Roz earns $18 for each hour she works. Which table represents the equation y = 18x?

OpenStudy (blackhawk92444):

just make table

OpenStudy (blackhawk92444):

are you just typing letters

OpenStudy (anonymous):

what look i just need help

OpenStudy (anonymous):

@fluttershyK

OpenStudy (jonw91):

y = 18x where x represents number of hours worked at $18 and hour. If I work 0 hours I get paid nothing, 0*18=0. If I work 1 hour I get paid $18, 1*18=18 If I work 2 hours I get paid $36, 2*18=36 If I work 3 hours I get paid $54, 3*18=54 If I turn that into a chart: y x 0 0 1 18 2 36 so on and so on.

OpenStudy (jonw91):

@fluttershyK - A is wrong. A shows starting at $18 and increasing by $1 or y=18+(x+1)

OpenStudy (anonymous):

ok

OpenStudy (fluttershyk):

oh yeah ok yep I was wrong im not good at math